OpenStudy (anonymous):

f(x)=([(x^4)/4] + [1/(8x^2)] on the interval 1,2 find the arc length. i need help on this please

OpenStudy (anonymous):

I am not sure (has been a long time). But the answer may be: \[ds = \sqrt{dx^{2}+dy^{2}}\] \[\frac{ds}{dx} = \sqrt{1+\left(\frac{dy^{2}}{dx^{2}}\right)}\] \[= \sqrt{1+\left(\frac{dy}{dx}\right)^{2}}\] \[\int f(x)\cdot ds = \int\sqrt{1+\left(\frac{df}{dx}\right)^{2}}dx\] \[f(x) = \frac{x^{4}}{4}+\frac{1}{8x^{2}}\] \[\frac{df}{dx} = x^{3}-\frac{1}{4x^{3}}\] \[1+\left(\frac{df}{dx}\right)^{2} = 1+x^{6}-2\cdot x^{3}\cdot\frac{1}{4x^{3}}+\frac{1}{16x^{6}}\] \[ = x^{6}+\frac{1}{2}+\frac{1}{16x^{6}}\] \[\sqrt{1+\left(\frac{df}{dx}\right)^{2}} = \sqrt{x^{6}+\frac{1}{2}+\frac{1}{16x^{6}}}\] \[ = \sqrt{\frac{16x^{12}+8x^{6}+1}{16x^{6}}}\] \[ = \frac{4x^{6}+1}{4x^{3}}\] \[ = 4x^{3}+4x^{-3}\] \[\int\sqrt{1+\left(\frac{df}{dx}\right)^{2}} = \frac{4x^{4}}{4}+4\cdot\left(-\frac{1}{2}\right)\cdot x^{-2}\] \[ = x^{4}-2x^{-2}\] \[\int_{1}^{2}\sqrt{1+\left(\frac{df}{dx}\right)^{2}} = 2^{4}-2\cdot2^{-2}-\left(1^{4}-2\cdot1^{-2}\right)\] \[ = 16-\frac{1}{4}-1+2\] \[ = 17.25\]